OpenStudy (anonymous):

A circle has its center at (1, 4) and a radius of 2 units. What is the equation of the circle? (1 point) Select one: a. (x + 2)2 + (y + 4)2 = 2 b. (x - 1)2 + (y - 4)2 = 4 c. (x + 1)2 + (y - 4)2 = 4 d. (x - 1)2 + (y - 4)2 = 2

jigglypuff314 (jigglypuff314):

equation of circle: (x-h)^2 + (y-k)^2 = r^2 when r = radius and (h, k) is the center so if the center is (1, 4) and radius is 2 plug it into the equation and see what you get :)

OpenStudy (anonymous):

so it would be answer b right?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

thank you

jigglypuff314 (jigglypuff314):

your welcome :D